What is the pH of the solution containing 0.20 M NH3 and 0.15 M NH4Cl?

1
Expert's answer
2022-05-05T04:21:03-0400

Kb (NH3) = 1.8•10-5

pKb= 4.74

pOH = pK- log (NH4Cl/NH3)

pOH = 4.74 - log(0.2/0.15)= 4.62

pH = 14 - pOH = 14 - 4.62 = 9.38